Manuel is 144 years old. Victoria is 38 years old. How many years until Manuel's age is 3 times Victoria's age?
Answer by fractalier(6550)

Call the number of years you are looking for, x.
Then we can write
144 + x = 3(38 + x)
144 + x = 114 + 3x
Now subtract x and subtract 114 from both sides and get
144 = 114 + 2x
30 = 2x
so that
x = 15 years in the future...